An organism has a lag period of 30 minutes and a generation time of 20 minutes. If the original concentration of a culture is 200 organisms/mL, what will be the concentration after 90 minutes?
Answer #1

Lag phase: Bacterial growth during this phase usually do not increase in number.Bacteria prepares for division and cell size is increased.

Log phase/exponential phase/generation time: Organism doubles in number in this time.

CALCULATIONS:

Given, lag period= 30 minutes.

Generation time= 20 minutes.

Total time given= 90 minutes.

Time left after lag time=(total time)-(lag time)=[(90min) -(30min)]=60 min

Number of generations organism will divide in this 60min=(Total time)/(generation time)= 3 generations

given initial concentration=200 organisms/mL.

organism doubles after each generation.so after first generation ,concentration=400organisms/ml

after second generation=800organisms/ml

after third generation=1600organisms/ml.

   therefore after 90 min ,total concentration of organisms=1600organisms/ml